# Electron-phonon coupling induced topological phase transition in an $\alpha$-$T_{3}$ Haldane-Holstein model

### v1

@@ Abstract
We present impelling evidence of topological phase transitions induced by electron-phonon (e-ph) coupling in an $\alpha$-$T_{3}$ Haldane-Holstein model that presents smooth tunability between graphene ($\alpha=0$) and a dice lattice $(\alpha=1)$. The e-ph coupling has been incorporated via the Lang-Firsov transformation which adequately captures the polaron physics in the high frequency (anti-adiabatic) regime, and yields an effective Hamiltonian of the system through zero phonon averaging at $T=0$. While exploring the signature of the phase transition driven by polaron and its interplay with the parameter $\alpha$, we identify two regions based on the values of $\alpha$, namely, the low to intermediate range $(0<\alpha\leq 0.6)$ and larger values of $\alpha~{}(0.6<\alpha<1)$ where the topological transitions show distinct behaviour. There exists a single critical e-ph coupling strength for the former, below which the system behaves as a topological insulator characterized by edge modes, finite Chern number, and Hall conductivity, with all of them vanishing above this value, and the system undergoes a spectral gap closing transition. Further, the critical coupling strength depends upon $\alpha$. For the latter case $(0.6<\alpha<1)$, the scenario is more interesting where there are two critical values of the e-ph coupling at which trivial-topological-trivial and topological-topological-trivial phase transitions occur for $\alpha$ in the range $[0.6:1]$. Our studies on e-ph coupling induced phase transitions show a significant difference with regard to the well-known unique transition occurring at $\alpha=0.5$ (or at $0.7$) in the absence of the e-ph coupling, and thus underscore the importance of interaction effects on the topological phase transitions.
@@ I Introduction
It has been proposed that the dice lattice can be realized in a cold-atom experimental set-up by three counter-propagating laser beams Bercioux et al. (2009). Furthermore, the realization of such lattice is proposed by growing a heterostructure consisting of a trilayer of cubic lattices, namely, SrTiO3/SrIrO3/SrTiO3 in the (111) direction Wang and Ran (2011). Interstingly, a smooth transformation from pseudospin $S=1/2$ state to pseudospin $S=1$ can be apprehended using a more generalized version of the $T_{3}$ system, known as the $\alpha-T_{3}$ lattice Raoux et al. (2014); Malcolm and Nicol (2015); Illes et al. (2015); Wang and Liu (2021), where $\alpha$ is the strength of the nearest neighbour (NN) hopping from the central atom (shown by the red line in Fig. 1). Therefore, an $\alpha$-$T_{3}$ lattice sets the two extreme limits, namely, $\alpha=0$ (graphene) and $\alpha=1$ (dice), between which a continuous control is possible by parametrizing $\alpha$ suitably through a variation of the Berry phase, proportional to $\alpha$. Raoux et al. Raoux et al. (2014) have explored the role of the Berry phase in $\alpha$-$T_{3}$ lattice and shown that the orbital susceptibility of the system undergoes a transition from a diamagnetic ($\alpha=0$) to a paramagnetic ($\alpha=1$) at a critical $\alpha_{c}=0.495$ while the Berry phase changes from $\pi$ (graphene) to 00 (dice). Due to the presence of the additional atom at the centre, the low-energy spectrum of the $\alpha$-$T_{3}$ lattice, governed by the tight-binding Dirac-Weyl Hamiltonian, embodies two dispersive bands that are linear in momenta and a dispersionless flat band at zero-energy. Malcolm et al. Malcolm and Nicol (2015) have shown that the bulk dispersion of a quantum well structure made of Hg1-xCdxTe can be linked to the low-lying dispersion of an $\alpha$-$T_{3}$ lattice with an effective $\alpha=1/\sqrt{3}$ at a critical Cd doping concentration, $x_{c}\approx 0.17$. In recent times, such $\alpha$-$T_{3}$ lattices with enlarged spin states ($S>1/2$) have offered a series of studies on equilibrium Raoux et al. (2014); Malcolm and Nicol (2015); Illes et al. (2015); Wang et al. (2020a); Wang and Liu (2021); Kovács et al. (2017); Islam and Dutta (2017); Illes and Nicol (2017); Biswas and Ghosh (2016, 2018); Singh and Sharma (2023a); Gorbar et al. (2019); Chen et al. (2019a); Illes and Nicol (2016); Chen et al. (2019b); Roslyak et al. (2021); Balassis et al. (2020); Wang et al. (2020b); Han and Lai (2022); Iurov et al. (2022, 2020); Singh and Sharma (2023b); Islam et al. (2023); Islam and Kapri (2022) and nonequilibrium Dey and Ghosh (2018, 2019); Mojarro et al. (2020); Niu and Wang (2022); Tamang et al. (2021); Tamang and Biswas (2023); Iurov et al. (2019) properties of $\alpha$-$T_{3}$ lattice. To highlight a few, people have investigated the role of Berry phase Raoux et al. (2014); Illes et al. (2015); Dey and Ghosh (2018); Iurov et al. (2019); Singh and Sharma (2023a), valley-polarized transport Islam and Dutta (2017); Niu and Wang (2022) that may be applicable to valleytronics, Klein tunneling Illes and Nicol (2017), optical conductivity Illes et al. (2015); Illes and Nicol (2016); Kovács et al. (2017); Chen et al. (2019a, b); Han and Lai (2022), magnetotransport properties, such as Shubnikov–de Hass oscillation and quantized Hall conductivity Biswas and Ghosh (2016); Illes et al. (2015); Wang et al. (2020a); Singh and Sharma (2023a), floquet dynamics Dey and Ghosh (2019); Tamang et al. (2021), and other topological signatures Mondal and Basu (2023); Bercioux et al. (2011); Wang and Liu (2021) in $\alpha-T_{3}$ lattices. As there was evidence in the past which demonstrated the topology of the multiband systems, such as, kagomé Tang et al. (2011); Liu et al. (2012); Trescher and Bergholtz (2012), Lieb Jaworowski et al. (2015), and dice Xu and Duan (2017); Vigh et al. (2013); Wang and Ran (2011); Bercioux et al. (2011); Dey et al. (2020); Mondal and Basu (2023) lattices could be turned into a Chern insulator by tuning the system parameters, recently, Dey et al. Dey and Ghosh (2019) have attempted to show the topological transition at a critical $\alpha_{c}=1/\sqrt{2}$ via irradiating an $\alpha$-$T_{3}$ lattice with a circularly polarized light.
@@
Figure 1: The schematic diagram of an $\alpha$-$T_{3}$ lattice is shown, where the red, purple, and green circles represent the sublattices A, B, and C sublattices, respectively. The NN hopping strength between A and B sublattices is $t$, while it is $t^{\prime}=\alpha t$ between B and C sublattices. The NNN hopping is $\lambda/\lambda^{\prime}e^{\phi_{ij/jk}}$ ($\lambda/\lambda^{\prime}e^{-\phi_{ij/jk}}$) for the clockwise (anticlockwise) direction. Here $\lambda^{\prime}=\alpha\lambda$.
II ${\boldsymbol{\alpha\--T_{3}}}$ lattice with electron-phonon interaction
@@
II.1 Haldane-Holstein model for ${\boldsymbol{\alpha\--T_{3}}}$ lattice
@@
$\mathcal{H}$ $=$ $\biggl{[}-t\sum_{\langle i,j\rangle}c^{\dagger}_{i}c_{j}-\alpha t\sum_{\langle j,k\rangle}c^{\dagger}_{j}c_{k}-\frac{i\lambda}{3\sqrt{3}}\sum_{\langle\langle i,j\rangle\rangle}e^{i\phi_{ij}}c^{\dagger}_{i}c_{j}$ (1)
$~{}-\frac{i\alpha\lambda}{3\sqrt{3}}\sum_{\langle\langle j,k\rangle\rangle}e^{i\phi_{jk}}c^{\dagger}_{j}c_{k}+H.C.\biggr{]}+M\sum_{i}c^{\dagger}_{i}S_{z}c_{i}$
@@ II.2 Polaronic Hamiltonian: Lang-Firsov approach
This is a coherent transformation of a displaced harmonic oscillator that eliminates the phonon degrees of freedom and transforms the Hamiltonian into that for an effective electronic system. We must specify that this unitary transformation works well in the high-frequency (non-adiabatic) regime, meaning the LO frequency of the phonons is much larger than the other electronic parameters of the system, i.e. when $\hbar\omega_{0}\gg t,t^{\prime},\lambda,\lambda^{\prime},M$ and $\lambda_{eph}$. The LFT transforms the total Hamiltonian (1) as (see Appendix 24 for the derivation)
$\tilde{\mathcal{H}}$ $=$ $-t\biggl{[}\sum_{\langle i,j\rangle}c^{\dagger}_{i}c_{j}e^{[X_{i}-X_{j}]}+\alpha\sum_{\langle j,k\rangle}c^{\dagger}_{j}c_{k}e^{[X_{j}-X_{k}]}\biggr{]}-\frac{i\lambda}{3\sqrt{3}}\biggl{[}\sum_{\langle\langle i,j\rangle\rangle}e^{\phi_{ij}}c^{\dagger}_{i}c_{j}e^{[X_{i}-X_{j}]}+\alpha\sum_{\langle\langle j,k\rangle\rangle}e^{\phi_{jk}}c^{\dagger}_{j}c_{k}e^{[X_{j}-X_{k}]}\biggr{]}$ (4)
@@
The $X$-terms in the exponent contain the phonon operators as:
$X_{i}=\lambda_{eph}(b_{i}-b^{\dagger}_{i}).$ (5)
At this stage, to eliminate the phonon degrees of freedom, one can obtain a zero-phonon average (at $T=0$), which reads as
@@
~$\tilde{\mathcal{H}}_{eff}$ $=$ $-\tilde{t}~{}\biggl{[}\sum_{\langle i,j\rangle}c^{\dagger}_{i}c_{j}+\alpha\sum_{\langle j,k\rangle}c^{\dagger}_{j}c_{k}\biggr{]}$ (7)
$~{}~{}-\frac{i\tilde{\lambda}}{3\sqrt{3}}\biggl{[}\sum_{\langle\langle i,j\rangle\rangle}e^{\phi_{ij}}c^{\dagger}_{i}c_{j}+\alpha\sum_{\langle\langle j,k\rangle\rangle}e^{\phi_{jk}}c^{\dagger}_{j}c_{k}\biggr{]}$
@@
II.3 The continuum ${\boldsymbol{\alpha\--T_{3}}}$-Holstein Hamiltonian
@@
~$\tilde{\mathcal{H}}(\bf{k})$ $=$ $-\tilde{t}~{}(h_{x}S_{x}+h_{y}S_{y})+\frac{2\tilde{\lambda}\mathfrak{Im}(f_{\textbf{k}})}{3\sqrt{3}~{}\cos\varphi}S_{zH}$ (9)
@@
~$\tilde{\mathcal{H}}(\bf{q})$ $=$ $\hbar\tilde{v_{f}}\begin{pmatrix}M-\lambda_{eph}^{2}\hbar\omega_{0}&(\nu q_{x}-iq_{y})~{}$cos$\varphi&0\\ (\nu q_{x}+iq_{y})~{}$cos$\varphi&-\lambda_{eph}^{2}\hbar\omega_{0}&(\nu q_{x}-iq_{y})~{}$sin$\varphi\\ 0&(\nu q_{x}+iq_{y})~{}$sin$\varphi&-M-\lambda_{eph}^{2}\hbar\omega_{0}\end{pmatrix}+\frac{\tilde{\lambda}\nu}{\cos\varphi}\begin{pmatrix}-$cos$\varphi&0&0\\ 0&$cos$\varphi-$sin$\varphi&0\\ 0&0&$sin$\varphi\end{pmatrix},$ (13)
@@
III Polaron induced topological features in an ${\boldsymbol{\alpha\--T_{3}}}$ lattice
@@ III.1 Bulk spectral properties
Figure 2: The bulk band structures with energy $E$ (in the units of $t$) of the bare Haldane model are shown along the dimensionless $k_{x}$ axis (at $k_{y}=0$) for various values of $\alpha$: $(a)$ $\alpha=0.4$, $(b)$ $\alpha=0.5$, and $(c)$ $\alpha=0.6$. Bands are no longer symmetric under the exchange of valleys (K and $\bf{K^{\prime}}$). The Haldane term taken as $\lambda=0.1t$.
In our study, all the energy parameters are taken in units of $t$, which is set to unity. Further, we fix $a=1$ (lattice constant) and $\hbar=1$, for convenience. Before delving into the specifics of the e-ph interaction, let us briefly explore the bare Haldane $\alpha$-$T_{3}$ lattice. In the absence of e-ph interaction and the mass term, and solely due to the breaking of the time-reversal symmetry by the Haldane term, the original zero-energy FB may get distorted. Additionally, the electronic band structure experiences valley splitting. Fig. 2 illustrates the low-energy bands for various values of $\alpha$ within the first BZ. The red, green, and blue colours denote the CB, the FB and the VBs, respectively. In three-band systems, there can be two distinct band gaps at the Dirac points: the gaps between (i) the CB and the distorted FB ($\Delta^{\textbf{K}/\textbf{K}^{\prime}}_{cf}$), and (ii) the distorted FB and VBs ($\Delta^{\textbf{K}/\textbf{K}^{\prime}}_{vf}$) at the K/$\textbf{K}/\textbf{K}^{\prime}$ points. The middle band exhibits no dispersion at $\alpha=0$ (not depicted here), but it gets more dispersive with increasing $\alpha$. We see a mild dispersive nature of the FB at $\alpha=0.4$ as shown in Fig. 2$(a)$. In Fig. 2$(b)$, where $\alpha=0.5$, the distorted FB now connects with the VB by closing the gap between them at the K valley, while in the other valley ($\textbf{K}^{\prime}$), the distorted FB connects to the CB. With further increase in $\alpha$, the gap reopens, as depicted in Fig. 2$(c)$ for $\alpha=0.6$. In case of $\alpha=1$ (not shown here), the spectral gap attains its maximum value, and the distorted FB regains its dispersionless behaviour. Notably, at $\alpha=0.5$ one finds that, $\Delta^{\textbf{K}}_{cf}\neq 0$, but $\Delta^{\textbf{K}}_{vf}=0$, and further $\Delta^{\textbf{K}^{\prime}}_{cf}=0$, whereas $\Delta^{\textbf{K}^{\prime}}_{vf}\neq 0$. These findings precisely correspond with the previously reported results concerning the $\alpha$-$T_{3}$ lattice  Wang and Liu (2021); Dey and Ghosh (2019).
@@
Figure 5: Plots of polaronic bulk band structure with energy $E$ (in the units of $t$) for $\alpha=0.7$ are shown along the dimensionless $k_{x}$ axis (at $k_{y}=0$) for $(a)$ $\lambda_{eph}<\lambda_{c_{1}}$ ($\lambda_{eph}=0.2$), $(b)$ $\lambda_{eph}=\lambda_{c_{1}}=0.28$, $(c)$ $\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$ ($\lambda_{eph}=0.35$), $(d)$ $\lambda_{eph}=\lambda_{c_{2}}=0.43$, and $(e)$ $\lambda_{eph}>\lambda_{c_{2}}$ ($\lambda_{eph}=0.6$). The parameters are taken as $\lambda=0.1t$ and $M=0.05t$. Further, $t$ and $\lambda$ values are modified as $\tilde{t}$ and $\tilde{\lambda}$ as mentioned in the text.
Figure 6: Plots of polaronic bulk band structure with energy $E$ (in the units of $t$) for $\alpha=0.8$ are shown along the dimensionless $k_{x}$ axis (at $k_{y}=0$) for $(a)$ $\lambda_{eph}=0$, $(b)$ $\lambda_{eph}<\lambda_{c_{1}}$ ($\lambda_{eph}=0.15$), $(c)$ $\lambda_{eph}=\lambda_{c_{1}}=0.2$, $(d)$ $\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$ ($\lambda_{eph}=0.35$), $(e)$ $\lambda_{eph}=\lambda_{c_{2}}=0.44$, and $(f)$ $\lambda_{eph}>\lambda_{c_{2}}$ ($\lambda_{eph}=0.6$). The parameters are taken as $\lambda=0.1t$ and $M=0.05t$. Further, $t$ and $\lambda$ values are modified as $\tilde{t}$ and $\tilde{\lambda}$ as mentioned in the text.

### v2

@@ Abstract
We present impelling evidence of topological phase transitions induced by electron-phonon (e-ph) coupling in an $\alpha$-$T_{3}$ Haldane-Holstein model that facilitates smooth tunability between graphene ($\alpha=0$) and a dice lattice $(\alpha=1)$. The e-ph coupling has been incorporated via the Lang-Firsov transformation which adequately captures the polaron physics in the high frequency (anti-adiabatic) regime, and yields an effective Hamiltonian through zero phonon averaging at $T=0$. While exploring the signature of phase transitions driven by polaron and its interplay with the parameter $\alpha$, we identify two regions based on the values of $\alpha$, namely, the low to intermediate range $(0<\alpha\leq 0.6)$ and larger values of $\alpha~{}(0.6<\alpha<1)$, where the topological transitions host distinct behaviour. There exists a single critical e-ph coupling strength for the former, below which the system behaves as a topological insulator characterized by edge modes, finite Chern number, and Hall conductivity, with all of them vanishing above this value, and the system undergoes a spectral gap closing transition. Further, the critical coupling strength depends upon $\alpha$. For the latter case $(0.6<\alpha<1)$, the scenario is more interesting where there are two critical values of the e-ph coupling at which trivial-topological-trivial and topological-topological-trivial phase transitions occur. Our study shows a significant difference with regard to the well-known unique transition occurring at $\alpha=0.5$ (or at $0.7$) in the absence of the e-ph coupling, and thus underscores the importance of interaction effects on topological phase transitions.
@@ I Introduction
It has been proposed that the dice lattice can be realized in a cold-atom experimental set-up by three counter-propagating laser beams Bercioux et al. (2009). Furthermore, the realization of such lattice is proposed by growing a heterostructure consisting of a trilayer of cubic lattices, namely, SrTiO3/SrIrO3/SrTiO3 in the (111) direction Wang and Ran (2011). Interstingly, a smooth transformation from pseudospin $S=1/2$ state to pseudospin $S=1$ can be apprehended using a more generalized version of the $T_{3}$ system, known as the $\alpha$-$T_{3}$ lattice Raoux et al. (2014); Malcolm and Nicol (2015); Illes et al. (2015); Wang and Liu (2021), where $\alpha$ is the strength of the nearest neighbour (NN) hopping from the central atom (shown by the red line in Fig. 1). Therefore, an $\alpha$-$T_{3}$ lattice sets the two extreme limits, namely, $\alpha=0$ (graphene) and $\alpha=1$ (dice), between which a continuous control is possible by parametrizing $\alpha$ suitably through a variation of the Berry phase, proportional to $\alpha$. Raoux et al. Raoux et al. (2014) have explored the role of the Berry phase in $\alpha$-$T_{3}$ lattice and shown that the orbital susceptibility of the system undergoes a transition from a diamagnetic ($\alpha=0$) to a paramagnetic ($\alpha=1$) at a critical $\alpha_{c}=0.495$ while the Berry phase changes from $\pi$ (graphene) to 00 (dice). Due to the presence of the additional atom at the centre, the low-energy spectrum of the $\alpha$-$T_{3}$ lattice, governed by the tight-binding Dirac-Weyl Hamiltonian, embodies two dispersive bands that are linear in momenta and a dispersionless flat band at zero-energy. Malcolm et al. Malcolm and Nicol (2015) have shown that the bulk dispersion of a quantum well structure made of Hg1-xCdxTe can be linked to the low-lying dispersion of an $\alpha$-$T_{3}$ lattice with an effective $\alpha=1/\sqrt{3}$ at a critical Cd doping concentration, $x_{c}\approx 0.17$. In recent times, such $\alpha$-$T_{3}$ lattices with enlarged spin states ($S>1/2$) have offered a series of studies on equilibrium Raoux et al. (2014); Malcolm and Nicol (2015); Illes et al. (2015); Wang et al. (2020a); Wang and Liu (2021); Kovács et al. (2017); Islam and Dutta (2017); Illes and Nicol (2017); Biswas and Ghosh (2016, 2018); Singh and Sharma (2023a); Gorbar et al. (2019); Chen et al. (2019a); Illes and Nicol (2016); Chen et al. (2019b); Roslyak et al. (2021); Balassis et al. (2020); Wang et al. (2020b); Han and Lai (2022); Sun et al. (2022); Iurov et al. (2022, 2020); Singh and Sharma (2023b); Islam et al. (2023); Islam and Kapri (2022) and nonequilibrium Dey and Ghosh (2018, 2019); Mojarro et al. (2020); Niu and Wang (2022); Tamang et al. (2021); Tamang and Biswas (2023); Iurov et al. (2019) properties. To highlight a few, people have investigated the role of Berry phase Raoux et al. (2014); Illes et al. (2015); Dey and Ghosh (2018); Iurov et al. (2019); Singh and Sharma (2023a), valley-polarized transport Islam and Dutta (2017); Niu and Wang (2022) that may be applicable to valleytronics, Klein tunneling Illes and Nicol (2017), optical conductivity Illes et al. (2015); Illes and Nicol (2016); Kovács et al. (2017); Chen et al. (2019a, b); Han and Lai (2022), magnetotransport properties, such as Shubnikov–de Hass oscillation and quantized Hall conductivity Biswas and Ghosh (2016); Illes et al. (2015); Wang et al. (2020a); Singh and Sharma (2023a), floquet dynamics Dey and Ghosh (2019); Tamang et al. (2021), and other topological signatures Mondal and Basu (2023); Bercioux et al. (2011); Wang and Liu (2021) in $\alpha$-$T_{3}$ lattices. As there was evidence in the past which demonstrated the topology of the multiband systems, such as, kagomé Tang et al. (2011); Liu et al. (2012); Trescher and Bergholtz (2012), Lieb Jaworowski et al. (2015), and dice Xu and Duan (2017); Vigh et al. (2013); Wang and Ran (2011); Bercioux et al. (2011); Dey et al. (2020); Mondal and Basu (2023) lattices could be turned into a Chern insulator by tuning the system parameters, recently, Dey et al. Dey and Ghosh (2019) have attempted to show the topological transition at a critical $\alpha_{c}=1/\sqrt{2}$ via irradiating an $\alpha$-$T_{3}$ lattice with a circularly polarized light.
@@
Figure 1: The schematic diagram of an $\alpha$-$T_{3}$ lattice is shown, where the red, purple, and green circles represent the sublattices A, B, and C sublattices, respectively. The NN hopping strength between A and B sublattices (solid black line) is $t$, while it is $t^{\prime}=\alpha t$ between A and C sublattices (solid red line). The NNN hopping between A-B-A (dashed red) or B-A-B (dashed purple) is $\lambda e^{i\phi_{ij}}$, while through C, it is $\lambda^{\prime}e^{i\phi_{jk}}$ between A-C-A (dashed red) and C-A-C (dashed green). Here $\lambda^{\prime}=\alpha\lambda$ and the phase $\phi_{ij}$ ($-\phi_{ij}$) denotes the clockwise (anticlockwise) direction.
II $\alpha$-$T_{3}$ lattice with electron-phonon interaction
@@
II.1 Haldane-Holstein model for $\alpha$-$T_{3}$ lattice
@@
$\mathcal{H}$ $=$ $\biggl{[}-t\sum_{\langle i,j\rangle}c^{\dagger}_{i}c_{j}-\alpha t\sum_{\langle j,k\rangle}c^{\dagger}_{j}c_{k}-\frac{\lambda}{3\sqrt{3}}\sum_{\langle\langle i,j\rangle\rangle}e^{i\phi_{ij}}c^{\dagger}_{i}c_{j}$ (1)
$~{}-\frac{\alpha\lambda}{3\sqrt{3}}\sum_{\langle\langle j,k\rangle\rangle}e^{i\phi_{jk}}c^{\dagger}_{j}c_{k}+H.C.\biggr{]}+M\sum_{i}c^{\dagger}_{i}S_{z}c_{i}$
@@ II.2 Polaronic Hamiltonian: Lang-Firsov approach
This is a coherent transformation of a displaced harmonic oscillator that eliminates the phonon degrees of freedom and transforms the Hamiltonian into that for an effective electronic system. We must specify that this unitary transformation works well in the high-frequency (non-adiabatic) regime, meaning the LO frequency of the phonons is much larger than the other electronic parameters of the system, i.e., when $\omega_{0}\gg t,t^{\prime},\lambda,\lambda^{\prime},M$ and $\lambda_{eph}$. The LFT transforms the total Hamiltonian (1) as (see Appendix 24 for the derivation)
$\tilde{\mathcal{H}}$ $=$ $-t\biggl{[}\sum_{\langle i,j\rangle}c^{\dagger}_{i}c_{j}e^{[X_{i}-X_{j}]}+\alpha\sum_{\langle j,k\rangle}c^{\dagger}_{j}c_{k}e^{[X_{j}-X_{k}]}\biggr{]}-\frac{\lambda}{3\sqrt{3}}\biggl{[}\sum_{\langle\langle i,j\rangle\rangle}e^{i\phi_{ij}}c^{\dagger}_{i}c_{j}e^{[X_{i}-X_{j}]}+\alpha\sum_{\langle\langle j,k\rangle\rangle}e^{i\phi_{jk}}c^{\dagger}_{j}c_{k}e^{[X_{j}-X_{k}]}\biggr{]}$ (4)
@@
The $X$-terms in the exponent contain the phonon operators as
$X_{i}=\lambda_{eph}(b^{\dagger}_{i}-b_{i}).$ (5)
At this stage, to eliminate the phonon degrees of freedom, one can obtain a zero-phonon average (at $T=0$), which reads for the exponents as
@@
~$\tilde{\mathcal{H}}_{eff}$ $=$ $\langle 0|\tilde{\mathcal{H}}|0\rangle=-\tilde{t}~{}\biggl{[}\sum_{\langle i,j\rangle}c^{\dagger}_{i}c_{j}+\alpha\sum_{\langle j,k\rangle}c^{\dagger}_{j}c_{k}\biggr{]}$ (7)
$~{}~{}-\frac{\tilde{\lambda}}{3\sqrt{3}}\biggl{[}\sum_{\langle\langle i,j\rangle\rangle}e^{i\phi_{ij}}c^{\dagger}_{i}c_{j}+\alpha\sum_{\langle\langle j,k\rangle\rangle}e^{i\phi_{jk}}c^{\dagger}_{j}c_{k}\biggr{]}$
@@
II.3 The continuum $\alpha$-$T_{3}$-Holstein Hamiltonian
@@
~$\tilde{\mathcal{H}}(\bf{k})$ $=$ $-\tilde{t}~{}(h_{x}S_{x}+h_{y}S_{y})-\frac{2\tilde{\lambda}\mathfrak{Im}(f_{\textbf{k}})}{3\sqrt{3}~{}\cos\varphi}S_{zH}$ (9)
@@
~$\tilde{\mathcal{H}}(\bf{q})$ $=$ $\hbar\tilde{v_{f}}\begin{pmatrix}M-\lambda_{eph}^{2}\hbar\omega_{0}&(\nu q_{x}-iq_{y})~{}$cos$\varphi&0\\ (\nu q_{x}+iq_{y})~{}$cos$\varphi&-\lambda_{eph}^{2}\hbar\omega_{0}&(\nu q_{x}-iq_{y})~{}$sin$\varphi\\ 0&(\nu q_{x}+iq_{y})~{}$sin$\varphi&-M-\lambda_{eph}^{2}\hbar\omega_{0}\end{pmatrix}-\frac{\tilde{\lambda}\nu}{\cos\varphi}\begin{pmatrix}-$cos$\varphi&0&0\\ 0&$cos$\varphi-$sin$\varphi&0\\ 0&0&$sin$\varphi\end{pmatrix},$ (13)
@@
III Polaron induced topological features in an $\alpha$-$T_{3}$ lattice
@@ III.1 Bulk spectral properties
Figure 2: The bulk band structures with energy $E$ (in the units of $t$) of the bare Haldane model are shown as a function of dimensionless momenta, $k_{x}$ (multiplied by the lattice constant) at $k_{y}=0$ for various values of $\alpha$: $(a)$ $\alpha=0.4$, $(b)$ $\alpha=0.5$, and $(c)$ $\alpha=0.6$. The red, green, and blue colours represent the CB, the FB and the VBs, respectively. Bands are no longer symmetric under the exchange of valleys (K and $\bf{K^{\prime}}$). The Haldane term is taken as $\lambda=0.1t$.
In our study, all the energy parameters are taken in units of $t$, which is set to unity. Further, we fix $a=1$ (lattice constant), $\phi_{ij}=\pi/2$ (the Haldane flux) and $\hbar=1$, for convenience.
Before delving into the specifics of the e-ph interaction, let us briefly explore the bare Haldane $\alpha$-$T_{3}$ lattice. In the absence of e-ph interaction and the mass term, and solely due to the breaking of the time-reversal symmetry by the Haldane term, the original zero-energy FB may get distorted. Additionally, the electronic band structure experiences valley splitting. Fig. 2 illustrates the low-energy bands for various values of $\alpha$ within the first BZ. The red, green, and blue colours denote the CB, the FB and the VBs, respectively. In three-band systems, there can be two distinct band gaps at the Dirac points: the gaps between (i) the CB and the distorted FB ($\Delta^{\textbf{K}/\textbf{K}^{\prime}}_{cf}$), and (ii) the distorted FB and VBs ($\Delta^{\textbf{K}/\textbf{K}^{\prime}}_{vf}$) at the K/$\textbf{K}/\textbf{K}^{\prime}$ points. The middle band exhibits no dispersion at $\alpha=0$ (not depicted here), but it gets more dispersive with increasing $\alpha$. We see a mild dispersive nature of the FB at $\alpha=0.4$ as shown in Fig. 2$(a)$. In Fig. 2$(b)$, where $\alpha=0.5$, the distorted FB now connects with the VB by closing the gap between them at the K valley, while in the other valley ($\textbf{K}^{\prime}$), the distorted FB connects to the CB. With further increase in $\alpha$, the gap re-opens, as depicted in Fig. 2$(c)$ for $\alpha=0.6$. In case of $\alpha=1$ (not shown here), the spectral gap attains its maximum value, and the distorted FB regains its dispersionless behaviour. Notably, at $\alpha=0.5$ one finds that, $\Delta^{\textbf{K}}_{cf}\neq 0$, but $\Delta^{\textbf{K}}_{vf}=0$, and further $\Delta^{\textbf{K}^{\prime}}_{cf}=0$, whereas $\Delta^{\textbf{K}^{\prime}}_{vf}\neq 0$. These findings precisely correspond with the previously reported results concerning the $\alpha$-$T_{3}$ lattice  Wang and Liu (2021); Dey and Ghosh (2019).
@@
Figure 5: Plots of polaronic bulk band structure with energy $E$ (in the units of $t$) for $\alpha=0.7$ are shown as a function of dimensionless momenta, $k_{x}$ (multiplied by the lattice constant) at $k_{y}=0$ for $(a)$ $\lambda_{eph}=0$, $(b)$ $\lambda_{eph}<\lambda_{c_{1}}$ ($\lambda_{eph}=0.2$), $(c)$ $\lambda_{eph}=\lambda_{c_{1}}=0.28$, $(d)$ $\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$ ($\lambda_{eph}=0.35$), $(e)$ $\lambda_{eph}=\lambda_{c_{2}}=0.43$, and $(f)$ $\lambda_{eph}>\lambda_{c_{2}}$ ($\lambda_{eph}=0.6$). The red, green, and blue colours represent the CB, the FB and the VBs, respectively. The parameters are taken as $\lambda=0.1t$ and $M=0.05t$. Further, $t$ and $\lambda$ values are modified as $\tilde{t}$ and $\tilde{\lambda}$ as mentioned in the text. The values of $\lambda_{c_{1}}$ and $\lambda_{c_{2}}$ are mentioned in Table 2.
Figure 6: Plots of polaronic bulk band structure with energy $E$ (in the units of $t$) for $\alpha=0.8$ are shown as a function of dimensionless momenta, $k_{x}$ (multiplied by the lattice constant) at $k_{y}=0$ for $(a)$ $\lambda_{eph}=0$, $(b)$ $\lambda_{eph}<\lambda_{c_{1}}$ ($\lambda_{eph}=0.15$), $(c)$ $\lambda_{eph}=\lambda_{c_{1}}=0.2$, $(d)$ $\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$ ($\lambda_{eph}=0.35$), $(e)$ $\lambda_{eph}=\lambda_{c_{2}}=0.44$, and $(f)$ $\lambda_{eph}>\lambda_{c_{2}}$ ($\lambda_{eph}=0.6$). The red, green, and blue colours represent the CB, the FB and the VBs, respectively. The parameters are taken as $\lambda=0.1t$ and $M=0.05t$. Further, $t$ and $\lambda$ values are modified as $\tilde{t}$ and $\tilde{\lambda}$ as mentioned in the text. The values of $\lambda_{c_{1}}$ and $\lambda_{c_{2}}$ are mentioned in Table 2.